Milk of magnesia, Mg(OH)2, is commonly used as an antacid and is available in both tablet and liquid forms.
c) You have been supplied with a concentrated oral suspension of milk of magnesia that has a concentration of 2400 mg/10mL. Explain how you would prepare 200.0 mL of a

600 mg/10mL solution of the antacid for use for a child.
1
Expert's answer
2020-05-01T14:41:27-0400

Let V(ml) be the volume of initial suspension. Then V will contain V/10*2400 = 240*V mg of milk of magnesia. We need 200/10*600 = 12000 mg of milk of magnesia.

240*V = 12000

V = 50 (ml)

So to prepare desired solution we need to take a 200 ml volumetric flask, add 50 ml of concentrated suspension, and add 150 ml (to the mark) of water.